# Black-hole spectroscopy from a giant quantum vortex

### v1

@@
Pietro Smaniotto pietro.smaniotto@nottingham.ac.uk School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Leonardo Solidoro School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Patrik Švančara School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Sam Patrick Department of Physics, King’s College London, University of London, Strand, London, WC2R 2LS, UK    Maurício Richartz Centro de Matemática, Computação e Cognição, Universidade Federal do ABC (UFABC), 09210-170 Santo André, São Paulo, Brazil    Carlo F. Barenghi School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, NE1 7RU, UK    Ruth Gregory Department of Physics, King’s College London, University of London, Strand, London, WC2R 2LS, UK Perimeter Institute, 31 Caroline Street North, Waterloo, ON, N2L 2Y5, Canada    Silke Weinfurtner silke.weinfurtner@nottingham.ac. School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK Perimeter Institute, 31 Caroline Street North, Waterloo, ON, N2L 2Y5, Canada Centre for the Mathematics and Theoretical Physics of Quantum Non-Equilibrium Systems, University of Nottingham, Nottingham, NG7 2RD, UK
@@
Black-hole spectroscopy [1] aims to infer physical properties of black holes by detecting the spectrum of quasinormal modes (QNMs) they emit while settling toward equilibrium [2, 3, 4]. Unlike normal modes, which are resonances of energy-conserving systems, QNMs are damped oscillations that arise when a system loses energy. The detection of the full spectrum is challenging due to their rapid decay, limiting observations only to the longest-lived mode [5, 6]. Theoretical [7] and numerical [8] studies suggest that environmental confinement due to surrounding plasma or dark matter can modify the QNM spectrum and enhance its detectability. Here we show that spatial confinement similarly affects the noise spectrum of nanometre-scale interface waves surrounding a giant quantum vortex in superfluid helium-4, an experimentally accessible quantum system that emulates dynamics in rotating curved spacetime [9]. In the available parameter space, we identify regimes that exhibit a multitude of QNMs whose real and imaginary frequencies are shifted with respect to those of the unbounded system, in agreement with theoretical predictions [10]. Our results highlight the versatility of quantum liquids as experimental platforms for validating theoretical toolkits in spectral stability, demonstrating their effectiveness in simulating both rotating curved spacetime and relevant environmental effects.
@@
Recent advances in the experimental control of classical and quantum fluids [11, 12] have enabled the simulation of complex and fundamental astrophysical phenomena that are difficult to study directly. One notable example is the propagation of waves in a moving fluid, which mirrors the dynamics of a scalar field in a curved spacetime [13], providing a laboratory platform to explore the physics of rotating black holes. The rotational nature of such spacetimes gives rise to nontrivial effects, such as frame dragging [14] and the formation of an ergosphere, a region where no object can remain at rest relative to infinity. The rich phenomenology of rotating curved spacetimes thus offers a valuable framework for testing general relativity and probing potential signatures of dark matter and modified gravity theories [15]. To date, several aspects of black hole physics have been successfully reproduced in fluid systems, including Hawking radiation [16, 17, 18] and rotational superradiance [19, 20, 21]. The emission of waves related to QNMs has been studied numerically in fluids [22, 23, 24, 25], Bose–Einstein condensates [26], exciton-polariton superfluids [27], and optical solitons [28], while experimental observations have relied on surface (gravity) waves in water [29]. These media are often dissipative, and their excitations have short lifetimes, justifying the treatment of these systems as effectively open. The absence of boundaries allows energy and information to escape to infinity.
Here, we examine the occurrence of QNMs in a newly developed gravity simulator [9] that employs surface waves (simply waves from here on) propagating on superfluid helium-4, a quantum liquid with negligible viscosity, approximately 100-times less than water [30]. Broadband fluctuations of the interface height (interface noise) are generated by minute mechanical oscillations of our experimental set-up and consist of a large number of superimposed waves. Due to the incredibly weak damping of waves in superfluid helium, the relatively small size of the system, and our ability to readout the interface noise over an extended period of time, the presence of boundaries cannot be ignored. Using a semi-analytical model [10], we have shown that spatial confinement modifies the resonance spectrum of gravitationally compact objects. The corresponding QNM frequencies shift, while their decay rates decrease, increasing the likelihood of simultaneously detecting multiple QNMs – including the longest-lived (fundamental) mode and its overtones, which experience greater damping [5]. Our experimental investigation confirms this prediction, revealing the enhanced excitation of multiple resonant modes whose frequency shifts match our model.
@@
Figure 1: Radial phase-space trajectories and scattering potentials. a Schematic top view of the experimental zone (one quadrant shown). The black region represents the drain opening ($5.0$ mm radius). The orange-shaded area marks the accessible field of view for interface wave detection, spanning radii from $8.2$ to $30.0$ mm (black dotted lines in later panels). The grey region denotes the glass wall at $37.3$ mm. b Azimuthal velocity component of the superfluid flow as a function of radius. Black points represent values averaged over $1.8$-mm radial intervals. The orange line fits Eq. (1), yielding circulation $C=(583\pm 3)$ mm2/s and a solid-body rotation frequency $\Omega=(0.10\pm 0.02)$ rad/s. Vertical error bars indicate one standard deviation. c Phase-space diagram for the $m=-12$ mode. The radial component of the wave vector $\bm{k}$ is denoted by $p$. Low-frequency modes (black and green arrowed lines) reflect on both sides of the vortex-induced effective potential, while high-frequency modes (yellow and orange lines) exceeding the potential’s maximum of $9.96$ Hz are allowed to propagate between the giant vortex core and the outer boundary (grey dashed line). d Effective scattering potential, evaluated from Eq. (4) at $p=0$, for four azimuthal modes (solid coloured lines). Maxima of the potential (coloured circles) correspond to light-ring frequencies. When circulation increases to $2C$ (dash-dotted coloured lines, maxima marked by crosses), the cavity between the light ring and the outer boundary (grey dashed line) shrinks in size and depth.
Our approach to explore these elusive excitations is based on a large draining vortex of superfluid helium in a cylindrical container [9]. The flow is driven by a spinning propeller acting as a centrifugal pump [31], establishing a steady recirculation loop of the superfluid between the outer glass wall and the central region, which contains a circular opening, as illustrated in Fig. 1a. The giant vortex manifests itself as a distinct, approximately $15$ mm deep and $5$ mm wide interface depression centred above the drain. Since the vortex strength (circulation) in superfluid helium is quantum-mechanically discretised [32], the core of the giant vortex underneath the interface consists of a dense, tangled array of $N_{C}=2\pi C/\kappa=36,\!600\pm 200$ singly-quantised vortices, each carrying a circulation quantum $\kappa=9.98\times 10^{-8}$ m2/s, yielding the circulation $C$ of the macroscopic flow field. As $N_{C}\gg 1$, the discrete nature of the vortices can be ignored, with the giant vortex core acting as a region of continuous vorticity [32]. The rotating, curved spacetime induced by the vortex is then probed by waves that interact with its flow field. Our primary measurable is therefore the spatially and time-resolved height of the superfluid interface, which we obtain by a unique implementation of synthetic Schlieren imaging [33, 34] in a low-temperature environment. This non-invasive optical technique captures small-amplitude interface height fluctuations within an annular field of view, as illustrated in Fig. 1a.
@@
Cylindrical symmetry and stationarity of the flow ensure the conservation of frequency $f$ and azimuthal number $m$, with the latter quantifying the symmetry of a wave along a circular path. We apply a discrete Fourier transform to the time and azimuthal coordinates to decompose the height fluctuation field into individual modes labelled by $m$ and $f$. Using the technique established in [9], we exploit the characteristic shape of wave spectra in Fourier space to determine the background velocity near the interface. We find that the radial velocity component is negligible in the fluid layer probed by the waves [35], while the azimuthal component exhibits a distinct radial dependence, shown by the black points in Fig. 1b. We verify that the flow field $\bm{v}$, given by
@@
is almost entirely irrotational, with the contribution from the small solid-body rotation term $\Omega r$ being subdominant with respect to the irrotational term $C/r$. The best fit of Eq. (1) to the experimental data is shown in Fig. 1b by an orange line.
@@
where $c$ represents the wave propagation speed. Although this metric description does not account for the role of dispersion in influencing the dynamics of high-frequency waves, previous studies have shown that curved spacetime phenomenology persists in more general regimes [19, 29, 36, 16, 18, 17]. To analyse the underlying wave dynamics, we employ the Wentzel–Kramers–Brillouin (WKB) approximation [37], which models interface fluctuations as a superposition of plane waves with a spatially varying wave vector $\bm{k}=p\,\hat{\bm{r}}+q\,\hat{\bm{\theta}}$ and fixed angular frequency $\omega=2\pi f$. The waves satisfy the dispersion relation [36],
$\left(\omega-\bm{v}\cdot\bm{k}\right)^{2}=F(k)\,,$ (3)
where $k\equiv||\bm{k}||$ and $F(k)$ is the function that characterises dispersive properties of the medium (see Methods). Flow symmetries constrain the azimuthal component of the wave vector to $q=m/r$, but the radial component $p$ is a free variable. Consequently, Eq. (3) admits propagating waves when $p$ is real valued. We can find these waves as intersections of a line of constant $\omega$ with one of the two branches of the dispersion relation,
$\omega_{D}^{\pm}(p,m,r)=\frac{mC}{r^{2}}+m\Omega\,\pm\sqrt{F(p,m,r)}\,.$ (4)
@@
For the flow parameters considered here, we find at most two propagating modes at each point, corresponding to radially ingoing ($p<0$) and outgoing ($p>0$) waves. In Fig. 1c, we show how the $p$ values of these two modes vary with radius for selected frequencies. The total wave vector defines the local wavelength, $2\pi/k$, which is on the order of millimetres near the outer boundary (grey dashed line) and decreases toward the centre.
@@
We identify two distinct behaviours. At low frequency, the ingoing wave starting near the boundary is reflected by the vortex, whereas, at higher frequency, the ingoing wave propagates all the way into the centre. We understand this behaviour as follows. For $p=0$, Eq. (4) exhibits a local minimum for $\omega_{D}^{+}$ and a maximum for $\omega_{D}^{-}$. The frequency range between these extrema does not contain real-valued solutions, instead admitting evanescent modes characterised by imaginary $p$. The curve $\omega_{D}^{+}(p=0,m,r)$ then separates the propagating and evanescent solutions, acting as an effective potential barrier.
@@
QNMs arise in scenarios where the effective potential features a local maximum [3]. In our system, this occurs only for waves counter-rotating with the vortex. In Fourier space, such waves are characterised by negative azimuthal numbers and, for selected values of $m$, the effective potential is shown in Fig. 1d as a function of radius. The location of the maximum is analogous to the light ring in black hole physics [38, 39]. It marks the unstable boundary between waves that fall into the black-hole horizon and those that escape to infinity, as shown in Fig. 1c, and is responsible for the shape of the accretion disk in images of real black holes [40]. In our case, waves near the light ring represent the lowest-energy excitations able to radiate energy from the vortex, and approximate the real part of the fundamental QNM frequency when the WKB approximation holds.
@@
A key factor shaping the QNM spectrum in our simulator is the presence of a boundary, which, in combination with the effective potential, forms a shallow cavity. The depth and spatial extent of this cavity are determined primarily by the circulation of the vortex. In Fig. 1d, effective potentials at double circulation values are shown by dash-dotted lines, illustrating how the maxima shift outward and the depth of the cavity decreases. This behaviour indicates that cavity effects can only be studied within a specific range of experimental parameters, notably constraining the feasible vortex circulations. We accessed this regime through a systematic exploration of the experimental parameter space, finding effective potentials suitable for investigating confinement effects across a range of superfluid temperatures and flow drive strengths. Here, we discuss a particular realisation at temperature $1.62$ K, with a driving propeller frequency of $1.67$ Hz. Under these conditions, we observe numerous resonances, which are related to the QNMs of an equivalent system, featuring the same effective potential but lacking the outer boundary.
@@
Figure 2: Black-hole spectroscopy. a, c Ensemble-averaged spectra of surface waves (black lines) with one standard deviation intervals (grey-shaded areas) for azimuthal numbers $m=-7$ and $-12$. Orange-shaded frames indicate the frequency range for the resonant mode search, delimited by the value of the effective potential at the outer boundary and the light-ring frequency incremented by 1 Hz. Green vertical lines mark the predicted resonances from panels b and d. b, d Maps of $-\log_{10}(\mathrm{Res})$ in the complex frequency plane, where $\mathrm{Res}$ is defined by Eq. (5). White dashed lines indicate the light-ring frequency. White symbols mark the resonances: circles for quasibound states (below the light-ring frequency) and crosses for other QNMs (above). White arrowed lines show resonance migration paths as the reflectivity of the outer boundary smoothly varies from 0 to 1. e Ensemble-averaged spectrogram of waves as a function of azimuthal number $m$. Positive (negative) $m$ corresponds to co-rotating (counter-rotating) waves. Bound states appear for $m>0$, while numerous QNMs emerge for $m<0$. Bright vertical stripes at $m=\pm 1$ are experimental artifacts. The orange dotted rectangle marks the region analysed in panel f. The colour bar follows a logarithmic scale. f Detailed view of the resonances from panel e. Each $m$-channel is ensemble-averaged and normalised by its mean amplitude. Low-frequency quasibound states (white circles) are prominent for $m\gtrsim-10$, while higher-frequency QNMs (white crosses) extend over a broader range of azimuthal numbers. The dashed white line marks the light-ring frequency. Dash-dotted white lines delineate the region of interest, consistent with the frequency extent in panels a and c.
@@
The interaction of a wave with the effective potential can be understood as a scattering process characterised by frequency-dependent reflection and transmission coefficients, respectively $R(\omega)$ and $T(\omega)$, which are the ratios of the reflected and transmitted wave amplitudes to the incident wave amplitude. The purely outgoing boundary condition, indicating the absence of the outer boundary, is achieved by setting the amplitude of the incoming wave to zero, and the corresponding QNM frequencies are identified as the poles of $R(\omega)$ and $T(\omega)$ in the complex frequency plane. For a reflective boundary at a finite radius, the amplitudes of the ingoing and outgoing waves must satisfy an appropriate boundary condition. We found the full-slip (Neumann) boundary condition to be consistent with experimental data. Resonant modes in such a system must satisfy $\mathrm{Res}(\omega)=0$, where
@@
The integration limits are set by the boundary position, $r_{B}=37.3$ mm, and the last scattering point, $r_{-}$, which is defined as follows. For frequencies below the light-ring frequency, $r_{-}$ is the crossing point of a line of constant $\omega$ with the effective potential. For higher frequencies, $r_{-}$ corresponds to the saddle point introduced in the WKB approximation [39] (see Methods). We emphasise that modes satisfying the resonance condition are directly linked to the QNMs of the corresponding open system. As shown in [10], one can gradually increase the boundary’s reflectivity and track how resonant modes shift from their original QNM frequencies to new resonances determined by evaluating Eq. (5). During this migration, the damping rate of resonant modes gets reduced, enhancing the visibility of multiple QNMs in spatially confined systems. This leads to the emergence of sharper, more distinct peaks in the frequency band just above the light-ring frequency.
@@
Our observations of the interface noise are presented in Fig. 2 and reveal a similarly rich spectral structure. As an example, we examine azimuthal numbers $m=-7$ and $-12$. Panels a and c show ensemble-averaged spectra (black lines) along with their standard deviations (grey-shaded areas). In the stationary flow studied here, steady excitations of specific modes enable us to average the noise spectra across overlapping time windows, reducing the contribution of random fluctuations. We perform radial averaging over a carefully selected region, excluding both the outer boundary and the giant vortex core. This approach allows us to focus on surface dynamics near the light ring (see Methods). The resulting spectra exhibit reduced amplitude at low frequencies, i.e. below the potential barrier. The first distinct amplitude peaks appear between approximately $5$ and $11$ Hz, around the respective light-ring frequencies, indicating the region where the least damped QNMs are predicted to be present.

### v2

@@
Pietro Smaniotto pietro.smaniotto@nottingham.ac.uk School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Leonardo Solidoro School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Patrik Švančara School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK    Sam Patrick Department of Physics, King’s College London, University of London, Strand, London, WC2R 2LS, UK    Maurício Richartz Centro de Matemática, Computação e Cognição, Universidade Federal do ABC (UFABC), 09210-170 Santo André, São Paulo, Brazil    Carlo F. Barenghi School of Mathematics, Statistics and Physics, Newcastle University, Newcastle upon Tyne, NE1 7RU, UK    Ruth Gregory Department of Physics, King’s College London, University of London, Strand, London, WC2R 2LS, UK Perimeter Institute, 31 Caroline Street North, Waterloo, ON, N2L 2Y5, Canada    Silke Weinfurtner silke.weinfurtner@nottingham.ac.uk School of Mathematical Sciences, University of Nottingham, University Park, Nottingham, NG7 2RD, UK Nottingham Centre of Gravity, University of Nottingham, University Park, Nottingham NG7 2RD, UK Perimeter Institute, 31 Caroline Street North, Waterloo, ON, N2L 2Y5, Canada Centre for the Mathematics and Theoretical Physics of Quantum Non-Equilibrium Systems, University of Nottingham, Nottingham, NG7 2RD, UK
@@
Black-hole spectroscopy [1] aims to infer physical properties of black holes by detecting the spectrum of quasinormal modes (QNMs) they emit while settling towards equilibrium [2, 3]. Unlike normal modes, which are resonances of energy-conserving systems, QNMs are damped oscillations arising when a system loses energy due to open boundaries or via dissipation [4]. The detection of the full QNM spectrum of black holes is challenging due to rapidly decaying amplitudes of these resonances, limiting observations only to the longest-lived mode [5, 6]. Theoretical [7] and numerical [8] studies suggest that environmental confinement due to surrounding plasma or dark matter modify the QNM spectrum. Here, we employ black-hole spectroscopy to show how spatial confinement similarly affects the spectrum of nanometre-scale interface waves surrounding a giant quantum vortex in superfluid helium-4, an experimentally accessible quantum system that emulates dynamics in rotating curved spacetime [9]. In the available parameter space, we observe regimes in which multiple QNMs emerge from the interface noise spectrum. In agreement with theoretical predictions [10], their real and imaginary frequencies are shifted with respect to those expected in the unbounded system. Our results demonstrate the critical role of spatial confinement in shaping the QNM spectrum, highlighting the importance of environmental effects on spectral stability of astrophysical compact objects.
@@
Recent advances in the experimental control of quantum fluids [11, 12, 13, 14, 15] have enabled the simulation of complex and fundamental astrophysical phenomena that are difficult to study directly. One notable example is the propagation of waves in a moving fluid that mirrors the dynamics of a scalar field in a curved spacetime [16], providing a laboratory platform to explore the physics of rotating black holes. The rotational nature of such spacetimes gives rise to nontrivial effects, such as the formation of an ergosphere, a region where no object can remain at rest relative to infinity. The rich phenomenology of rotating curved spacetimes thus offers a valuable framework for testing general relativity and probing potential signatures of dark matter and modified gravity theories [17]. To date, several aspects of black-hole physics have been successfully reproduced in fluid systems, including Hawking radiation [18, 19, 20] and rotational superradiance [21, 22, 23]. The emission of waves related to QNMs has been studied numerically in ideal fluids [24, 25, 26], Bose–Einstein condensates [27], exciton-polariton superfluids [28], and optical solitons [29]. In these models, the boundary conditions are open, permitting energy and information to escape to infinity. Experimental observations of QNMs emitted by a draining vortex in a water tank [30] likewise result in an effectively open system, because the waves are affected by linear dissipation and decay too quickly to reach the system’s boundary.
We examine the occurrence of QNMs in a newly developed gravity simulator [9] that employs superfluid helium-4, a quantum liquid with a viscosity approximately two orders of magnitude lower than that of water [31]. The effective rotating curved spacetime is induced by a giant quantum vortex that is stabilised by draining the superfluid through an opening in the centre of a cylindrical container (see Methods for further details). Mechanical vibrations of this cryogenic set-up induce interface noise, i.e. broadband fluctuations of the superfluid interface composed of many superimposed surface waves. We capture these fluctuations and hence probe the effective curved spacetime by uniquely implementing synthetic Schlieren imaging in a quantum liquid [32, 33], enabling non-invasive, space- and time-resolved detection of surface waves. Considering the incredibly weak damping of superfluid waves, the relatively small size of our system, and the ability to read out interface fluctuations over an extended period of time, boundary effects cannot be ignored. Using a toy-model of a spatially confined black-hole spacetime [10], it has been shown that QNM signatures persist in the simulated spectrum, with their frequencies shifting along the real axis while their damping rates decrease but do not vanish. This reduction in damping enhances the prospect of detecting multiple QNMs simultaneously, including the longest-lived fundamental mode and its strongly damped overtones. Here, we confirm this prediction experimentally and infer the effective boundary condition governing the wave dynamics at the system’s outer boundary, extending black-hole spectroscopy to the study of specific confinement effects.
@@
Figure 1: Radial phase-space trajectories and scattering potentials. a Schematic top view of the experimental zone (one quadrant shown). The black region represents the drain opening ($5.0$ mm radius). The orange-shaded area marks the accessible field of view for interface-wave detection, spanning radii from $8.2$ to $30.0$ mm (grey dotted lines in later panels). The grey region denotes the glass wall at $37.3$ mm. b Velocity components of the superfluid flow as a function of radius. Black (grey) points represent azimuthal (radial) velocity values averaged over $1.8$-mm radial intervals. The orange line fits Eq. (1), yielding circulation $C=(583\pm 3)$ mm2/s and a solid-body rotation frequency $\Omega=(0.10\pm 0.02)$ rad/s. Vertical error bars indicate one standard deviation. c Phase-space diagram for the $m=-12$ mode, calculated for the flow parameters extracted from panel b. The radial component of the wave vector $\bm{k}$ is denoted by $p$. Modes with frequencies below $9.96$ Hz reflect on both sides of the vortex-induced effective potential plotted in panel d, while higher-frequency modes are allowed to propagate between the giant vortex core and the outer boundary (grey dashed line). d The effective scattering potential, evaluated from Eq. (4) at $p=0$, for four azimuthal modes (solid coloured lines). Maxima of the potential (coloured circles) correspond to light-ring frequencies. If we were to consider double circulation $2C$ (dash-dotted coloured lines, maxima marked by crosses), the cavity between the light ring and the outer boundary (grey dashed line) would reduce in both size and depth.
@@
Fig. 1a outlines the geometry of our experiment. The giant quantum vortex is located above the central opening (black region) and manifests itself as a distinct, approximately $15$ mm deep and $5$ mm wide depression of the superfluid interface. Interface height fluctuations are resolved within an $21.8$-mm wide annular region (orange) that is surrounded by an outer boundary (grey). Cylindrical symmetry and stationarity of the flow ensure the conservation of frequency $f$ and azimuthal number $m$ of surface waves, with the latter quantifying the wave’s periodicity along a circular path. We apply a discrete Fourier transform to the time and azimuthal coordinates to decompose the interface noise into individual modes labelled by $m$ and $f$. Using the technique established in [9], we exploit the characteristic shape of the noise spectra in Fourier space to determine the background velocity near the interface. As shown in Fig. 1b, the radial velocity component (grey points) is negligible in the fluid layer probed by the waves [34], while the azimuthal component (black points) exhibits a distinct radial dependence. The overall flow field can be written as,
@@
with the dominant contribution from the irrotational term $C/r$, and a smaller (less than $10\%$ for $r<24$ mm) solid-body rotation term $\Omega r$. The best fit of Eq. (1) is shown in Fig. 1b by an orange line.
@@
where $c$ represents the wave propagation speed. Although this metric does not account how dispersion affects high-frequency waves, previous studies have shown that curved-spacetime phenomenology persists in such regimes [21, 30, 35, 18]. To analyse the underlying wave dynamics, we employ the Wentzel–Kramers–Brillouin (WKB) approximation [36], which models interface fluctuations as a superposition of plane waves with a spatially varying wave vector $\bm{k}=p\,\hat{\bm{r}}+(m/r)\,\hat{\bm{\theta}}$ and fixed angular frequency $\omega=2\pi f$. The waves satisfy the dispersion relation [35],
$\left(\omega-\bm{v}\cdot\bm{k}\right)^{2}=F(\bm{k})\,,$ (3)
where $F(\bm{k})$ is the function that characterises dispersive properties of the medium (see Methods). The radial wave vector $p(r)$ is obtained by solving the dispersion relation at each radius $r$. Equivalently, we can recast Eq. (3) as $\omega=\omega_{D}^{\pm}$ where,
$\omega_{D}^{\pm}(p,m,r)=\frac{mC}{r^{2}}+m\Omega\,\pm\sqrt{F(p,m,r)}\,,$ (4)
@@
and search for the intersections of fixed frequency lines with the curves $\omega_{D}^{\pm}$. When these intersections occur for $p\in\mathbb{R}$, the dispersion relation has propagating wave solutions. Given the flow parameters considered here, we find at most two propagating modes for each frequency, corresponding to radially ingoing ($p<0$) and outgoing ($p>0$) waves. In Fig. 1c, we compute the $p$ values of these two modes as a function of radius. The total wave vector defines the local wavelength, $2\pi/k$, which is on the order of millimetres near the outer boundary (grey dashed line) and decreases towards the centre.
@@
We identify two distinct behaviours. At low frequencies, the ingoing wave starting near the boundary is reflected by the vortex, whereas, at higher frequencies, the ingoing wave propagates all the way into the centre. We understand this behaviour as follows. For $p=0$, Eq. (4) exhibits a local minimum for $\omega_{D}^{+}$ and a maximum for $\omega_{D}^{-}$. The frequency range between these extrema does not contain real-valued solutions, instead admitting evanescent modes characterised by imaginary value of $p$. The curve $\omega_{D}^{+}(p=0,m,r)$ then separates the propagating and evanescent solutions, acting as an effective potential barrier.
@@
QNMs arise in scenarios where the effective potential features a local maximum [4]. In our system, this occurs only for waves counter-rotating with the vortex, which are characterised by negative azimuthal numbers. The corresponding effective potential is shown as a function of radius in Fig. 1d. The location of the maximum is analogous to the light ring in black-hole physics [37, 38]. It marks the unstable boundary between waves that fall into the black-hole horizon and those that escape to infinity, as shown in Fig. 1c, and is responsible for the shape of the accretion disk in images of real black holes [39]. In our case, waves near the light ring represent the lowest-energy excitations able to radiate energy from the vortex, and approximate the real part of the fundamental QNM frequency when the WKB approximation holds.
@@
A key factor shaping the QNM spectrum in our simulator is the presence of a boundary, which, in combination with the effective potential, forms a shallow cavity. The depth and spatial extent of this cavity are determined primarily by the circulation of the vortex. In Fig. 1d, effective potentials at double circulation values are shown by dash-dotted lines, illustrating how the maxima shift outward and depths of the cavities decrease. This behaviour indicates that cavity effects can only be studied within a specific range of experimental parameters, notably constraining the feasible vortex circulations. We accessed this regime through a systematic exploration of the experimental parameter space, finding effective potentials suitable for investigating confinement effects across a range of superfluid temperatures and flow drive strengths. Here, we discuss a particular realisation at temperature $1.62$ K, with a driving propeller frequency of $1.67$ Hz. Under these conditions, we observe numerous resonances, which are directly related to the QNMs of the same system, but in the absence of reflection at the outer boundary.
@@
Figure 2: Black-hole spectroscopy. a, c Ensemble-averaged spectra of surface waves (black lines) with one standard deviation intervals (grey-shaded areas) for azimuthal numbers $m=-4$ and $-16$. Orange-shaded frames indicate the frequency range for the resonant mode search, delimited by the value of the effective potential at the outer boundary and the light-ring frequency incremented by 1 Hz. Green vertical bands are centred at the predicted resonances from panels b and d. The width of each band reflects the statistical error in the estimation of $C$ and $\Omega$. Red ticks on the upper axis mark harmonics of the driving frequency. b, d Maps of $-\log_{10}(\mathrm{Res})$ in the complex frequency plane, where $\mathrm{Res}$ is defined by Eq. (5). White dashed lines indicate the light-ring frequency. White symbols mark the resonances: circles for quasibound states and crosses for other QNMs. Typical error of pinpointing these frequencies related to uncertainties in $C$ and $\Omega$ is estimated to be approx. $0.05$ Hz and $0.01$ Hz, for the real and imaginary part, respectively. White arrowed lines show resonance migration paths as the reflectivity of the outer boundary smoothly increases from 0 to 1. e Ensemble-averaged spectrogram of waves as a function of the azimuthal number $m$. Positive (negative) $m$ corresponds to co-rotating (counter-rotating) waves. Bound states appear for $m>0$, while numerous QNMs emerge for $m<0$. Bright vertical stripes at $m=\pm 1$ are experimental artifacts. The orange dotted rectangle marks the region analysed in panel f. The colour bar follows a logarithmic scale. f Detailed view of the resonances from panel e. Each $m$-channel is ensemble-averaged and normalised by its mean amplitude. Low-frequency quasibound states (white circles) are prominent for $m\lesssim-8$, while higher-frequency QNMs (white crosses) extend over a broader range of azimuthal numbers. The dashed white line marks the light-ring frequency. Dash-dotted white lines delineate the region of interest, consistent with the frequency extent in panels b and d.
@@
The interaction of a wave with the effective potential can be understood as a scattering process characterised by frequency-dependent reflection and transmission coefficients, respectively $R(\omega)$ and $T(\omega)$, which are the ratios of the reflected and transmitted wave amplitudes to the incident wave amplitude. The purely outgoing boundary condition, indicating the absence of the outer boundary, is achieved by setting the amplitude of the incoming wave to zero, and the corresponding QNM frequencies are identified as the poles of $R(\omega)$ and $T(\omega)$ in the complex frequency plane. For a reflective boundary at a finite radius, the amplitudes of the ingoing and outgoing waves must satisfy different constraints. Boundary conditions, which are consistent with experimental data, are full-slip (Neumann) at the external boundary and fully absorbing at the vortex core. The full-slip condition is suitable for inviscid waves interacting with a wall, while the fully absorbing condition aligns with the presence of a one-directional membrane, such as the event horizon considered in black-hole spectroscopy.
Resonant modes in such a system must satisfy $\mathrm{Res}(\omega)=0$, where
@@
The integration limits are set by the boundary position, $r_{B}=37.3$ mm, and the last scattering point, $r_{-}$, which is defined as follows. For frequencies below the light-ring frequency, $r_{-}$ is the crossing point of a line of constant $\omega$ with the effective potential. For higher frequencies, $r_{-}$ corresponds to the saddle point introduced in the WKB approximation [38] (see Methods). We emphasise that modes satisfying the resonance condition are directly linked to the QNMs of the corresponding open system. As shown in [10], one can gradually increase the boundary’s reflectivity and track how resonant modes shift from their original QNM frequencies to new resonances determined by evaluating Eq. (5). During this migration, the damping rate of resonant modes reduces, enhancing the visibility of multiple QNMs in spatially confined systems. This leads to the emergence of sharper, more distinct peaks in the frequency band just above the light-ring frequency.
@@
Our observations of the interface noise are presented in Fig. 2 and reveal a similarly rich spectral structure. As an example, we examine azimuthal numbers $m=-4$ and $-16$. Panels a and c show ensemble-averaged spectra (black lines) along with their standard deviations (grey-shaded areas). In the stationary flow studied here, steady excitations of specific modes enable us to average the noise spectra across overlapping time windows, reducing the contribution of random fluctuations. We perform radial averaging over a region excluding both the outer boundary and the giant vortex core. This approach allows us to focus on surface dynamics near the light ring (see Methods). Systematic spectral peaks across all azimuthal numbers appear at multiples of the drive frequency of $1.67$ Hz and are marked by red axis ticks. Excluding these excitations, the spectra feature reduced amplitude at frequencies below the effective potential barrier. First distinct spectral peaks emerge near the respective light-ring frequencies ($6.57$ Hz and $11.38$ Hz). These frequencies, together with the values of the effective potential at the boundary, define the (orange-shaded) regions, where the least damped QNMs are expected to occur.